OpenStudy (anonymous):

simplify (8 plus 1/x) divided by (5/x plus 4)

OpenStudy (anonymous):

\[\large 8+\frac{1}{x} \over \large \frac{5}{x}+4\]\[{\large 8+\frac{1}{x} \over \large \frac{5}{x}+4 }\times \frac{x}{x}=\frac{8x+1}{4x+5}\]hmm what next.. lol